# Approximation of multipartite quantum states: revised version with new applications

### v1

@@ 5.1 Simplified definition of $E_{R,\pi}$
5.2 Finite-dimensional approximation of $E_{R}$ and $E_{R}^{\infty}$ and its use
@@
Many results describing properties of the relative entropy of $\pi$-entanglement and its regularization in finite-dimensional multipartite quantum systems remain valid in the infinite-dimensional case under some additional conditions. A simple way to establish the validity of these results is to use the approximation technique based on Theorem 4 and the results in [15].
@@
We illustrate this approach restricting attention to the relative entropy of entanglement $E_{R}$ and its regularization $E_{R}^{\infty}$ defined by formulae (36) and (39) in which $\pi$ is the finest partition $\left\{\{1\},\ldots,\{n\}\right\}$ (in this case $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A_{1}...A_{n}})$ coincides with the convex set $\mathfrak{S}_{\rm sep}(\mathcal{H}_{A_{1}...A_{n}})$ of all fully separable states).
@@
Proposition 4.  Let $\rho$ be a state from the set $\,\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$ defined in (40). Let $\rho_{k}=Q_{k}\rho\hskip 1.0ptQ_{k}[\mathrm{Tr}Q_{k}\rho\hskip 1.0pt]^{-1}$, $Q_{k}=P_{k}^{1}\otimes...\otimes P_{k}^{n}$, where $\{P_{k}^{1}\}\subset\mathfrak{B}(\mathcal{H}_{A_{1}})$,…, $\{P_{k}^{n}\}\subset\mathfrak{B}(\mathcal{H}_{A_{n}})$ are arbitrary sequences of projectors strongly converging to the unit operators $I_{A_{1}}$,…,$I_{A_{n}}$ correspondingly. Then
@@
To guarantee the validity of (50) with $E_{R}^{*}=E_{R}$ it suffices to require that $\rho$ is any state in $\,\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ having at least $\,n-1$ marginal states with finite entropy.
Proof. Assume that $\rho$ is any state in $\,\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ having at least $\,n-1$ marginal states with finite entropy. This assumptions holds for any state in $\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$.
By using the monotonicity of the quantum mutual information (QMI) under local quantum operations it is easy to show that
@@
Since $c_{k}\to 0$ as $k\to+\infty$, this inequality and the lower semicontinuity of QMI imply that
@@
The finiteness of $I(A_{1}\!:...:\!A_{n})_{\rho}$ follows from the upper bound (4).
By Proposition 3 in [30] the limit relation (52) shows that
If $E_{R}^{*}=E_{R}$ then (50) follows from (52) by Proposition 2 in [15]. If $E_{R}^{*}=E_{R}^{\infty}$ and $\rho$ is a state in $\,\mathfrak{S}_{*}(\mathcal{H}_{A_{1}...A_{n}})$ then (50) follows from (51) by Theorem 4B in Section 4. $\Box$
Below we consider examples of using Proposition 5.2.
@@
Proposition 5.2 allows us to show that the inequalities in (53) remain valid for any state $\rho$ of a bipartite infinite-dimensional system provided that the state $\rho_{A_{i}}$ has theFA-property [^15] This assumption implies, by Theorem 1 in [32], the finiteness of the extended conditional entropy $S(A_{i}|A_{j})_{\rho}$ defined in (3). and $S(\cdot|\cdot)$ is the extended conditional entropy defined in (3). Indeed, it suffices to show that
@@
where $P_{k}^{s}$ is the spectral projector of $\rho_{A_{s}}$ corresponding to its $r$ maximal eigenvalues, $s=1,2$. This can be done by using Theorem 3B, since the function $\rho\mapsto S(A_{1}|A_{2})_{\rho}$ belongs to the class $L_{2}^{1}(2,1)\subset II_{2}^{1}$ (see Section 2.3).
@@
Proposition 5.2 allows us to show that all the inequalities in (54) with $E_{R}^{*}=E_{R}$ (resp. with $E_{R}^{*}=E_{R}^{\infty}$) remain valid for any pure state $\rho$ of a tripartite infinite-dimensional system provided that any two of the states $\rho_{A_{1}}$, $\rho_{A_{2}}$ and $\rho_{A_{3}}$ have finite entropy (resp. have the FA-property). [^16] This assumption and the purity of the state $\rho$ imply, by Theorem 1 in [32], the finiteness of all the entropies $S(\rho_{A_{i}A_{j}})$ involved in (54). Indeed, it suffices to show that
$S(\rho_{A_{1}A_{2}})=\lim_{k\to+\infty}S([\rho_{k}]_{A_{1}A_{2}}),\quad\rho_{k}=Q_{k}\rho\hskip 1.0ptQ_{k}[\mathrm{Tr}Q_{k}\rho\hskip 1.0pt]^{-1},\quad Q_{k}=P_{k}^{1}\otimes P_{k}^{2}\otimes P_{k}^{3},$
@@
where $P_{k}^{s}$ is the spectral projector of $\rho_{A_{s}}$ corresponding to its $r$ maximal eigenvalues, $s=1,2,3$. This can be done by noting that $S(\rho_{A_{1}A_{2}})=S(\rho_{A_{3}})$ and $S([\rho_{k}]_{A_{1}A_{2}})=S([\rho_{k}]_{A_{3}})$ as the states $\rho$ and $\rho_{k}$ are pure. Since $c_{k}[\rho_{k}]_{A_{3}}\leq\rho_{A_{3}}$ for each $k$ for some positive number $c_{k}$ tending to $1$ as $k\to+\infty$, the required limit relation can be proved easily by using concavity and lower semicontinuity of the entropy.
5.3 On the energy-constrained versions of $E_{R,\pi}$

### v2

@@ 4.1.2 Simplified definition of $E_{R,\pi}$
4.1.3 Continuity and convexity properties of $E_{R,\pi}^{\infty}$ in infinite-dimensions
It is shown in [6] that the relative entropy of $\pi$-entanglement $E_{R,\pi}$ is a lower semicontinuous function on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{m}})$ for any set of partitions $\pi$. Then a criterion of local continuity of $E_{R,\pi}$ is obtained in [7], which allows us to derive several verifiable conditions of local continuity of $E_{R,\pi}$ for any set partitions $\pi$, in particular, to prove that continuity of $E_{R,\pi}$ on a subset of $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ follows from continuity of the quantum mutual information (QMI) on this subset (this is important as there are easy verifiable conditions for local continuity of the QMI [37, Section 4.2]).
Unfortunately, the approach used in [6, 7] cannot be applied to analysis of the regularized relative entropy of $\pi$-entanglement $E_{R,\pi}^{\infty}$ well defined by the expression (67) for any set $\pi$ consisting of a single partition. Analytical properties of the function $E^{\infty}_{R,\pi}$ can be explored by using the results from Section 3, since this function belongs to the class $II_{n}^{n-1}$ (in terms of Section 2.3) for any $\pi$ consisting of a single partition (this is mentioned in Section 4.1.1).
The continuity properties of $E_{R,\pi}^{\infty}$ are described in the following
Proposition 5.  Let $\,\widetilde{\mathfrak{S}}^{n-1}_{II}(\mathcal{H}_{A_{1}...A_{n}})$ be the subset of $\,\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ defined in (68). Let $\pi$ be the set consisting of a single partition of $\,\{1,2,..,n\}$.
A) The function $E^{\infty}_{R,\pi}$ is finite and lower semicontinuous on the set $\,\widetilde{\mathfrak{S}}^{n-1}_{II}(\mathcal{H}_{A_{1}...A_{n}})$. Moreover,
$\liminf_{k\to+\infty}E^{\infty}_{R,\pi}(\rho_{k})\geq E^{\infty}_{R,\pi}(\rho_{0})$ (71)
for any sequence $\{\rho_{k}\}\subset\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ converging to a state $\rho_{0}\in\widetilde{\mathfrak{S}}^{n-1}_{II}(\mathcal{H}_{A_{1}...A_{n}})$.
B) Let $\{\rho_{k}\}$ be an arbitrary sequence of states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ converging to a state $\rho_{0}\in\widetilde{\mathfrak{S}}^{n-1}_{II}(\mathcal{H}_{A_{1}...A_{n}})$. The limit relation
holds provided one of the following conditions is valid: [^19] The finiteness of $\,E_{R,\pi}(\rho_{0})$ and $\,I(A_{1}\!:...:\!A_{n})_{\rho_{0}}$ follows from the condition $\rho_{0}\in\widetilde{\mathfrak{S}}^{n-1}_{II}(\mathcal{H}_{A_{1}...A_{n}})$.
  1. $\rm a)$
    $\lim_{k\to+\infty}D(\rho_{k}\|\hskip 1.0pt\omega_{k})=D(\rho_{0}\|\hskip 1.0pt\omega_{0})<+\infty;$
  2. $\rm b)$
    there exists
    $\lim_{k\to+\infty}E_{R,\pi}(\rho_{k})=E_{R,\pi}(\rho_{0})<+\infty;$
  3. $\rm c)$
    there exists
    $\lim_{k\to+\infty}I(A_{1}\!:...:\!A_{n})_{\rho_{k}}=I(A_{1}\!:...:\!A_{n})_{\rho_{0}}<+\infty.$
Remark 12. The conditions $\rm a)$,$\,\rm b)$ and $\rm c)$ in Proposition 4.1.3B are related as follows: Theorem 1 in [7] states that $\,\rm a)\Rightarrow b)$ in general and that $\,\rm a)\Leftrightarrow b)$ if the state $\rho_{0}$ is faithful; Proposition 2 in [7] states that $\,\rm c)\Rightarrow b)$.
Easy verifiable conditions for the validity of $\,\rm c)$ are presented in [37, Section 4.2].
Proof. A) It is mentioned at the end of Section 4.1.1 that the function $E^{\infty}_{R,\pi}$ belongs to the class $II_{n}^{n-1}$ in the settings $A^{\prime}_{1}=A_{s(1)}$,…,$A^{\prime}_{n}=A_{s(n)}$ for any permutation $\{s(i)\}_{i=1}^{n}$ of $[n]$. So, this claim can be derived from claim $\rm(i)$ of the $II_{n}^{m}$-part of Corollary 3.2 in Section 3.2 by applying the arguments in Remark 4.1.1. It suffices only to note that the function $E^{\infty}_{R,\pi}$ does not increase under action of local channels.
B) By Remark 4.1.3 it suffices to prove that condition $\rm b)$ implies (72). Let $p\in\mathbb{N}$ be arbitrary. By Corollary 22 in [6] the cone generated by the set $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A^{p}_{1}|...|A^{p}_{n}})$ is a weak* closed subset of $\mathfrak{T}_{+}(\mathcal{H}_{A^{p}_{1}|...|A^{p}_{n}})$. So, since the state $\omega^{\prime}\otimes\omega^{\prime\prime}$ lies in $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A^{p}_{1}|...|A^{p}_{n}})$ for any states $\omega^{\prime}\in\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A^{t}_{1}|...|A^{t}_{n}})$ and $\omega^{\prime\prime}\in\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A^{p-t}_{1}|...|A^{p-t}_{n}})$, $t=1,2,...,p-1$ (because $\pi$ consists of a single partition), the condition $\rm b)$ and the repeated use of Corollary 2G in [7] allow us to show that
It follows that
Remark 13. The proof of Proposition 4.1.3B shows that the limit relation (73) holds for an arbitrary sequence $\{\rho_{k}\}$ of states in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ converging to any state $\rho_{0}$ in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ provided that one of the conditions $\rm a)$,$\,\rm b)$ and $\rm c)$ in Proposition 4.1.3B is valid.
The convexity of the function $E_{R,\pi}^{\infty}$ (defined by the expression (67) for any set $\pi$ consisting of a single partition) follows from the convexity and subadditivity of $E_{R,\pi}$ by Proposition 13 in [49] (the proof remains valid in the infinite-dimensional settings). Since the function $\rho\mapsto E_{R,\pi}(\rho)$ is lower semicontinuous on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ by Corollary 23 in [6], the function $\rho\mapsto k^{-1}E_{R,\pi}(\rho^{\otimes k})$ has the same property for any $k$. So, it follows from the definition (67) that $E_{R,\pi}^{\infty}$ is a measurable function on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ (w.r.t. the Borel $\sigma$-algebra in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$).
Unfortunately, the convexity and measurability of $E_{R,\pi}^{\infty}$ does not imply that
$\,E_{R,\pi}^{\infty}(\bar{\rho}(\mu))\leq\int_{\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})}\,E_{R,\pi}^{\infty}(\rho)\mu(d\rho).$ (74)
for any Borel probability measure $\mu$ on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$, where $\bar{\rho}(\mu)$ is the barycenter of $\mu$ defined in (4). This inequality could be deduced from a similar inequality for the function $\rho\mapsto k^{-1}E_{R,\pi}(\rho^{\otimes k})$ provided that the convexity of the latter function is proved. But the arguments from the proof of Proposition 13 in [49] (implying the convexity of $E_{R,\pi}^{\infty}$) do not imply the convexity of the function $\rho\mapsto k^{-1}E_{R,\pi}(\rho^{\otimes k})$ for a given finite $k$. Nevertheless, the validity of (74) under a certain condition on $\bar{\rho}(\mu)$ can be proved by using the results from Section 3.2. Since $E_{R,\pi}^{\infty}\in II_{n}^{n-1}$, the $II_{n}^{m}$-part of Corollary 3.2 implies the following
[…]
Note that the inequality (74) with $E_{R,\pi}^{\infty}$ replaced by $E_{R,\pi}$ (where $\pi$ is any set partitions) holds for any Borel probability measure $\mu$ on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ because the nonnegative convex function $E_{R,\pi}$ is lower semicontinuous on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ by Corollary 23 in [6] (the Jensen inequality (2) holds for any convex lower semicontinuous function $f$ on $\mathfrak{S}(\mathcal{H})$ and any Borel probability measure $\mu$ on $\mathfrak{S}(\mathcal{H})$, see, f.i., Proposition A-2 in [27]).
Remark 14. The claims of Propositions 4.1.3 and 4.1.3 in the bipartite case $n=2$ are also valid for the regularization $E_{R,PPT}^{\infty}$ of the relative entropy distance to the PPT states [50, 51]. This can be shown by repeating the arguments from the proofs of these claims, since the cone generated by the set of PPT states is weak* closed (by Lemma 24 in [6]) and the function $E_{R,PPT}^{\infty}$ belongs to the class $II^{1}_{2}$ (a faithful $II_{2}^{1}$-type continuity bound for $E_{R,PPT}^{\infty}$ is presented in Proposition 36 in [6]).
4.1.4 Finite-dimensional approximation of $E_{R}$ and $E_{R}^{\infty}$ and its use
@@
Many results describing properties of the relative entropy of $\pi$-entanglement and its regularization in finite-dimensional multipartite quantum systems remain valid in the infinite-dimensional case under some additional conditions. A simple way to establish the validity of these results is to use the approximation technique based on the continuity conditions in [7] and Proposition 4.1.3 in Section 4.1.3.
@@
We illustrate this approach restricting attention to the relative entropy of entanglement $E_{R}$ and its regularization $E_{R}^{\infty}$ defined by formulae (64) and (67) in which $\pi$ is a singleton set consisting of the finest partition $\left\{\{1\},\ldots,\{n\}\right\}$ (in this case $\mathfrak{S}_{\rm sep}^{\pi}(\mathcal{H}_{A_{1}...A_{n}})$ coincides with the convex set $\mathfrak{S}_{\rm sep}(\mathcal{H}_{A_{1}...A_{n}})$ of all fully separable states).
@@
Proposition 7.  Let $\rho$ be a state from the set $\,\widetilde{\mathfrak{S}}^{n-1}_{II}(\mathcal{H}_{A_{1}...A_{n}})$ defined in (68). Let $\rho_{k}=Q_{k}\rho\hskip 1.0ptQ_{k}[\mathrm{Tr}Q_{k}\rho\hskip 1.0pt]^{-1}$, $Q_{k}=P_{k}^{1}\otimes...\otimes P_{k}^{n}$, where $\{P_{k}^{1}\}\subset\mathfrak{B}(\mathcal{H}_{A_{1}})$,…, $\{P_{k}^{n}\}\subset\mathfrak{B}(\mathcal{H}_{A_{n}})$ are arbitrary sequences of projectors strongly converging to the unit operators $I_{A_{1}}$,…,$I_{A_{n}}$ correspondingly. Then
@@
To guarantee the validity of (75) with $E_{R}^{*}=E_{R}$ it suffices to require that $\rho$ is any state in $\,\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ such that $E_{R}(\rho)<+\infty$.
Proof. Assume first that $\rho$ is any state in $\,\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ such that $E_{R}(\rho)<+\infty$. By the monotonicity of the relative entropy of entanglement under local channels this implies $E_{R}(\rho_{A_{1}...A_{m}})<+\infty$. Consider the sequence of states
$\hat{\rho}_{k}=\hat{Q}_{k}\rho\hskip 1.0pt\hat{Q}_{k}[\mathrm{Tr}\hat{Q}_{k}\rho\hskip 1.0pt]^{-1},\quad\hat{Q}_{k}=P_{k}^{1}\otimes...\otimes P_{k}^{m}\otimes I_{A_{m+1}}\otimes...\otimes I_{A_{n}},$
tending to the state $\rho$. By the monotonicity of the relative entropy of entanglement under selective LOCC we have
$\hat{c}_{k}E_{R}([\hat{\rho}_{k}]_{A_{1}...A_{m}})\leq E_{R}(\rho_{A_{1}...A_{m}})<+\infty,\quad\hat{c}_{k}=[\mathrm{Tr}\hat{Q}_{k}\rho\hskip 1.0pt].$
Thus, by using the lower semicontinuity of $E_{R}$ (Corollary 20 in [6]) we obtain
$\lim_{k\to+\infty}E_{R}([\hat{\rho}_{k}]_{A_{1}...A_{m}})=E_{R}(\rho_{A_{1}...A_{m}})<+\infty.$
Since $c_{k}[\rho_{k}]_{A_{1}...A_{m}}\leq\hat{c}_{k}[\hat{\rho}_{k}]_{A_{1}...A_{m}}$ for all $k$, where $c_{k}=\mathrm{Tr}Q_{k}\rho$, and $c_{k},\hat{c}_{k}\to 1$ as $k\to+\infty$, this implies (75) with $E_{R}^{*}=E_{R}$ by the dominated convergence condition for $E_{R}$ [7, Proposition 2].
Assume now that $\rho$ is any state in $\,\widetilde{\mathfrak{S}}^{n-1}_{II}(\mathcal{H}_{A_{1}...A_{n}})$. Then $E_{R}(\rho_{A_{1}...A_{m}})<+\infty$ by inequality (66) because any state having the FA-property has finite entropy. So, it suffices to prove (75) with $E_{R}^{*}=E_{R}^{\infty}$. Since the finiteness of $I(A_{1}\!:...:\!A_{n})_{\rho}$ follows from the upper bound (9), this can be done by applying Proposition 4.1.3B in Section 4.1.3 with the help of Lemma 4.1.4 below. $\Box$
Lemma 5. Let $\rho$ be a state in $\,\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ such that $\,I(A_{1}\!:...:\!A_{n})_{\rho}<+\infty\,$ and $\{\rho_{k}\}$ be the sequence defined by the rule described in Proposition 4.1.4 via arbitrary sequences $\{P_{k}^{1}\}\subset\mathfrak{B}(\mathcal{H}_{A_{1}})$,…, $\{P_{k}^{n}\}\subset\mathfrak{B}(\mathcal{H}_{A_{n}})$ of projectors strongly converging to the unit operators $I_{A_{1}}$,…,$I_{A_{n}}$ correspondingly. Then
Proof. By using the monotonicity of the quantum mutual information (QMI) under local quantum operations it is easy to show that
@@
Since $c_{k}\to 1$ as $k\to+\infty$, it follows from this inequality and the lower semicontinuity of QMI that
@@
By the second claim of Proposition 4.6 in [37] the limit relation (77) imply (76). $\Box$
Below we consider examples of using Proposition 4.1.4.
@@
Proposition 4.1.4 allows us to show that the inequalities in (78) remain valid for any state $\rho$ of a bipartite infinite-dimensional quantum system provided that the state $\rho_{A_{i}}$ has finite entropy and $S(\cdot|\cdot)$ is the extended conditional entropy defined in (7). Indeed, by the additivity of the entropy and the relative entropy it suffices to show that (78) holds with $E_{R}^{\infty}(\rho)$ replaced by $E_{R}(\rho)$. Due to the inequality (66) this can be done by applying the last claim of Proposition 4.1.4. By symmetry we have only to show that
@@
where $P_{k}^{s}$ is the spectral projector of $\rho_{A_{s}}$ corresponding to its $k$ maximal eigenvalues, $s=1,2$. This can be done by using Theorem 3.1A, since the function $\rho\mapsto S(A_{1}|A_{2})_{\rho}$ belongs to the class $I_{2}^{1}$ (a faithful $I^{1}_{2}$-type continuity bound for this function is obtained by Winter in [23]).
@@
Proposition 4.1.4 allows us to show that
  • all the inequalities in (79) with $E_{R}^{*}=E_{R}$ remain valid (with possible value +$\infty$+ ∞ in one or both sides) for any pure state $\rho$ of a tripartite infinite-dimensional system;
@@
  • all the inequalities in (79) with with $E_{R}^{*}=E_{R}^{\infty}$ remain valid for any pure state $\rho$ of a tripartite infinite-dimensional system provided that any two of the states $\rho_{A_{1}}$, $\rho_{A_{2}}$ and $\rho_{A_{3}}$ have the FA-property. [^20] This assumption and the purity of the state $\rho$ imply, by Theorem 1 in [4] and inequality (66), the finiteness of all terms in (79).
Indeed, note first that in the case $E_{R}(\rho)=+\infty$ all the inequalities in (79) with $E_{R}^{*}=E_{R}$ hold trivially. Thus, to derive both claims from Proposition 4.1.4 it suffices (by symmetry) to show that
$\lim_{k\to+\infty}S([\rho_{k}]_{A_{1}A_{2}})=S(\rho_{A_{1}A_{2}})\leq+\infty,\quad\rho_{k}=Q_{k}\rho\hskip 1.0ptQ_{k}[\mathrm{Tr}Q_{k}\rho\hskip 1.0pt]^{-1},\quad Q_{k}=P_{k}^{1}\otimes P_{k}^{2}\otimes P_{k}^{3},$
@@
for any pure state $\rho$ of $A_{1}A_{2}A_{3}$, where $P_{k}^{s}$ is the spectral projector of $\rho_{A_{s}}$ corresponding to its $k$ maximal eigenvalues, $s=1,2,3$. This can be done by noting that $S(\rho_{A_{1}A_{2}})=S(\rho_{A_{3}})$ and $S([\rho_{k}]_{A_{1}A_{2}})=S([\rho_{k}]_{A_{3}})$ as the states $\rho$ and $\rho_{k}$ are pure. Since $c_{k}[\rho_{k}]_{A_{3}}\leq\rho_{A_{3}}$ for each $k$ for some positive number $c_{k}$ tending to $1$ as $k\to+\infty$, the required limit relation can be proved easily by using concavity and lower semicontinuity of the entropy.
Example 10. If $ABC$ is a finite-dimensional tripartite quantum system then
$I(A\!:\!B|C)_{\rho}\geq E^{\infty}_{R}(\rho_{A|BC})-E^{\infty}_{R}(\rho_{A|C})\quad\forall\rho\in\mathfrak{S}(\mathcal{H}_{ABC}),$ (80)
where $I(A\!:\!B|C)_{\rho}=S(\rho_{AC})+S(\rho_{BC})-S(\rho_{ABC})-S(\rho_{C})\,$ is the quantum conditional mutual information and $E^{\infty}_{R}(\rho_{X|YZ})$ denotes the regularization of the bipartite relative entropy of entanglement of a state $\rho_{XYZ}$ w.r.t. the partition $X|YZ$ [35, Lemma 1]. The proof of this result is based on the state redistribution protocol proposed in [53, 54]. So, if we want to upgrade this proof to the case of infinite-dimensional tripartite quantum system $ABC$ it is necessary to upgrade the results from [53, 54] to this case. It obviously requires serious technical efforts (to take the possibility of infinite values of the marginal entropies of a state into account, the necessity to use of extended definition of $I(A\!:\!B|C)_{\rho}$, etc.).
Proposition 4.1.4 allows us to directly derive an infinite-dimensional generalization of Lemma 1 in [35] from its original version. Namely, it allows us to prove that inequality (80) is valid for a state $\rho$ of infinite-dimensional tripartite quantum system $ABC$ provided that the FA-property holds either for the state $\rho_{A}$ (case $(a)$) or for both states $\rho_{B}$ and $\rho_{C}$ (case $(b)$) and
$I(A\!:\!B|C)_{\rho}=\left\{\begin{array}[]{l}I(A\!:\!BC)_{\rho}-I(A\!:\!C)_{\rho}\;\;\textit{in case }\;(a)\\ I(B\!:\!AC)_{\rho}-I(B\!:\!C)_{\rho}\;\;\textit{in case }\;(b).\end{array}\right.$ (81)
Since $\,\rho_{X}\in\mathfrak{S}_{\rm\textsf{FA}\!}(\mathcal{H}_{X})\,\Rightarrow S(\rho_{X})<+\infty\,\Rightarrow I(X\!:\!Y)_{\rho}<+\infty\,$ by the Theorem in [4] and due to (9), all the expressions in (81) are well defined. If all the marginal entropies of a state $\rho$ are finite then these expressions coincide with the above finite-dimensional formula for $I(A\!:\!B|C)_{\rho}$. They also agree with the extension of $I(A\!:\!B|C)_{\rho}$ to the set of all states of an infinite-dimensional tripartite quantum system $ABC$ [16, Theorem 2].
To prove the above generalization of Lemma 1 in [35] it suffices to take arbitrary sequences $\{P_{k}^{X}\}\subset\mathfrak{B}(\mathcal{H}_{X})$ of finite-rank projectors strongly converging to the unit operators $I_{X}$, $X=A,B,C$. Theorem 2 in [16] and Proposition 4.1.4 imply, respectively, that
$\lim_{k\to+\infty}I(A\!:\!B|C)_{\rho_{k}}=I(A\!:\!B|C)_{\rho}<+\infty,$
and
$\lim_{k\to+\infty}E_{R}^{\infty}([\rho_{k}]_{A|X})=E_{R}^{\infty}(\rho_{A|X})<+\infty,\quad X=B,BC$
in both cases $(a)$ and $(b)$. In case $(b)$ this is true, since in this case $\rho_{BC}\in\mathfrak{S}_{\rm\textsf{FA}\!}(\mathcal{H}_{BC})$ by Proposition 1 in [4]. By Lemma 1 in [35] the inequality (80) holds for the state $\rho_{k}\doteq Q_{k}\rho\hskip 1.0ptQ_{k}[\mathrm{Tr}Q_{k}\rho\hskip 1.0pt]^{-1}$, $Q_{k}=P_{k}^{A}\otimes P_{k}^{B}\otimes P_{k}^{C}$, for any $k$. So, the above limit relations imply that (80) holds for the state $\rho$.
4.1.5 On the energy-constrained versions of $E_{R,\pi}$